# A Structured Neural ODE Approach for Real Time Evaluation of AC Losses in 3D Superconducting Tapes

### v1

@@ Abstract
Efficient modeling of High Temperature Superconductors (HTS) is crucial for real-time quench monitoring; however, full-order electromagnetic simulations remain prohibitively costly due to the strong nonlinearities. Conventional reduced-order methods, such as the Proper Orthogonal Decomposition (POD) and Discrete Empirical Interpolation Method (DEIM), alleviate this cost but are limited by intrusive implementation and by the need for many interpolation points. This work investigates reduced-order strategies for Integral Equation Method (IEM) of HTS systems. We present the first application of PODDEIM to IEM-based HTS models, and introduce a Structured Neural Ordinary Differential Equation (Neural ODE) approach that learns nonlinear dynamics directly in the reduced space. Benchmark results show that the Neural ODE outperforms PODDEIM in both efficiency and accuracy, highlighting its potential for real-time superconducting simulations.
@@ 1 Introduction
Although HTS offer the major advantage of operating at significantly higher temperatures than conventional Low Temperature Superconductors (LTS), monitoring the quench phenomenon remains a critical challenge. This difficulty arises from the relatively slow reduction of the critical current $I_{c}$ with increasing temperature [6, 7]. The sudden transition from the superconducting to the normal resistive state generates localized hot spots through the Joule effect, which, if not properly controlled, can lead to irreversible damage of the conductor. Reliable quench detection and mitigation are therefore essential for the safe and continuous operation of future fusion demonstrator power plants, such as the European DEMO [8].
@@
Modeling quench phenomena requires solving multiphysics problems that couple electromagnetics (EM), thermal behavior, and fluid dynamics [9]. While lumped-element models have been proposed, capturing the underlying physics with sufficient accuracy typically demands finite element analysis. For superconducting structures, the EM problem can be addressed using several numerical formulations, most notably the Finite Element Method (FEM) with magnetic field ($H$) formulations [10], or the TTAA approach [11]. An alternative is the IEM, which is traditionally employed in open-domain problems such as antenna radiation, but has also proven effective for computing AC losses in superconductors [12, 13].
@@
Due to the strong nonlinearities in the resistivity model, both FEM and IEM simulations require substantial computational resources to capture the transient, nonlinear dynamics of realistic 3D superconducting structures. This high computational cost renders them impractical for real-time monitoring. To address this limitation, several studies have explored the use of surrogate models to accelerate the EM solution, particularly in the context of AC loss estimation [14, 15]. A first attempt to apply POD to HTS systems within the IEM framework was reported in [16]. However, in this work no hyperreduction strategy was adopted, as the authors argued that the sparsity of the operator $\mathbf{R}$ kept the assembly cost manageable. This is true for relatively small problems, but as the mesh dimension increases, the assembly quickly becomes considerably more demanding. In such cases, the Empirical Interpolation Method (EIM) [17] and its discrete counterpart [18] can be applied to alleviate this cost, and these techniques have already been employed successfully in electromagnetics problems [19]. Nevertheless, the DEIM suffers from two critical drawbacks in the present context: the strong nonlinearity and spatial locality of superconducting material laws require a large number of interpolation points, which reduces efficiency, and its intrusive nature demands full access to the high-fidelity model assembly routines, which limits practical applicability. Furthermore, since the governing equations are highly nonlinear, the reduced-order model must still be solved with an implicit time-integration scheme, which entails a large number of Newton iterations and repeated Jacobian evaluations. As a result, even with DEIM, the simulation cost remains significant.
@@
These challenges have motivated a growing interest in Machine Learning (ML)-based reduction strategies, which are typically non-intrusive and data-driven. Here, we provide only a brief overview of the main approaches, without aiming to be exhaustive; several recent reviews give a comprehensive account of this rapidly evolving area [20, 21, 22].
@@
Neural Operators such as DeepONet [23], the Fourier Neural Operator [24], and graph neural operator variants [25] aim to learn mappings from function spaces, rather than between finite-dimensional vectors. This formulation enables them to generalize across families of input functions, achieving mesh-independent predictions and fast inference once trained. Their flexibility and non-intrusivity make them particularly attractive for parametric problems and large-scale simulations. However, they also present drawbacks: training requires substantial data, the networks can be very large, and their ability to handle complex geometries is limited. Graph-based neural operators alleviate this to some extent, but at the price of predicting the full-order field directly, which implies very large network outputs and nontrivial computational costs.
@@
Physics-Informed Neural Networks (PINNs) [26] incorporate the residual of the physical equation into the training loss, enabling the solution of forward and inverse problems even when data are scarce. PINNs are especially powerful for inverse problems and for settings where measurements are sparse. However, they tend to struggle with complex geometries, and their formulation is not well-suited for input-driven or control problems: a PINN is typically trained for a fixed set of inputs or excitations, and adapting it to new ones generally requires retraining the network.
@@
Latent Dynamics modeling frameworks [29, 30] take yet another route. The central idea is that the solution trajectories of high-dimensional problems evolve on a low-dimensional manifold, which can be identified either with linear projections (e.g. POD) or with nonlinear encoders (e.g., autoencoders [31, 32]). Once the manifold is fixed, the latent dynamics can be learned in different ways. Sparse Identification of Nonlinear Dynamics (SINDy) [33] discovers sparse governing equations from a library of candidate functions, offering interpretability and parsimony. Neural Ordinary Differential Equations (Neural ODEs) [34] instead parametrize continuous-time dynamics through a trainable vector field.
@@
In this work, we build on the latent dynamics paradigm and propose a novel neural architecture for nonlinear electromagnetic modeling that operates directly in the reduced-order space. The starting point is a discretized integral equation model of HTS conductors, which is compressed through POD-based dimensionality reduction. Instead of relying on intrusive hyper-reduction techniques such as DEIM, we design a structured Neural ODE that learns only the nonlinear component of the reduced dynamics, while the linear part is inherited directly from the full-order model. In this way, physical knowledge is embedded into the reduced-order evolution, ensuring parsimony, interpretability, and efficiency. The approach is validated on a HTS CORC® cable case study and benchmarked against the standard PODDEIM method, demonstrating the potential of physics-informed latent dynamics modeling for challenging electromagnetic simulation. Results demonstrate improved accuracy and significantly lower computational cost, highlighting the potential of the proposed strategy for real-time, industrially relevant superconducting applications.
@@
  • We introduce and validate a Structured Neural ODE approach that directly learns the nonlinear dynamics in the reduced-order space, offering a non-intrusive alternative to traditional hyperreduction techniques.
@@ 2 Model Description
Figure 1: Resulting 2D mesh of the CORC® cable under analysis.
@@
We investigate the current density distribution induced by the magnetic field generated by nearby AC cables. Specifically, the magnetic field is assumed to be oriented in the $y$-direction, produced by the current flowing through an AC cable. This external field is modeled as:
@@
where $B_{0}$ is the magnetic field amplitude, which depends on both the AC current amplitude and the relative distance between the AC cable and the CORC® under study, and $f=50$ Hz is the frequency. According to Faraday’s law of induction, this time-varying magnetic field induces an electric field:
@@ 3 $J-\varphi_{e}$ Formulation
Although this formulation leads to non-sparse matrices, it is advantageous because only the regions carrying current—namely, the superconducting domains—need to be meshed. This avoids the need to mesh the surrounding air, unlike the $H$ and $H-\varphi_{e}$ formulations, which require solving for air elements as well. As a result, the overall model size is significantly reduced.
@@ 4.1 Structure-Preserving POD
    Using Full Order Model (FOM) simulations, snapshots of the current and potential are collected:
@@
    where $n_{e}$, $n_{f}$, and $n_{s}$ are the number of edges, faces, and snapshot instances, respectively.
@@
    $\mathbf{V}_{i,r}\in\mathbb{R}^{n_{e}\times r_{i}},\quad\mathbf{V}_{\varphi,r}\in\mathbb{R}^{n_{f}\times r_{\varphi}};$ (14)
@@ 4.3 POD-DEIM ROM Simulation
$\begin{bmatrix}\Delta t\mathbf{R}_{r}\big(\mathbf{i}_{r}^{(n+1)}\big)+\mathbf{L}_{r}&\Delta t\mathbf{G}_{r}^{T}\\ \Delta t\mathbf{G}_{r}&\mathbf{0}\end{bmatrix}\begin{bmatrix}\mathbf{i}_{r}^{(n+1)}\\ \mathbf{\varphi}_{r}^{(n+1)}\end{bmatrix}=\begin{bmatrix}\mathbf{L}_{r}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}\end{bmatrix}\begin{bmatrix}\mathbf{i}_{r}^{(n)}\\ \mathbf{\varphi}_{r}^{(n)}\end{bmatrix}+\begin{bmatrix}\Delta t\mathbf{e}_{s,r}^{(n+1)}\\ \mathbf{0}\end{bmatrix}$, (23)
with $\mathbf{R}_{r}$ is assembled using only the rows selected by the DEIM procedure. This discretization leads to a nonlinear system, since $\mathbf{R}_{r}$ depends nonlinearly on $\mathbf{i}_{r}^{(n+1)}$. Solving it requires a Newton solver, which in turn involves repeated Jacobian evaluations. This process is computationally expensive. To reduce the cost, we can, in principle, introduce an approximation: we evaluate $\mathbf{R}_{r}$ at the current from the previous time step $\mathbf{i}_{r}^{(n)}$, rather than at the unknown current $\mathbf{i}_{r}^{(n+1)}$:
$\begin{bmatrix}\Delta t\mathbf{R}_{r}\big(\mathbf{i}_{r}^{(n)}\big)+\mathbf{L}_{r}&\Delta t\mathbf{G}_{r}^{T}\\ \Delta t\mathbf{G}_{r}&\mathbf{0}\end{bmatrix}\begin{bmatrix}\mathbf{i}_{r}^{(n+1)}\\ \mathbf{\varphi}_{r}^{(n+1)}\end{bmatrix}=\begin{bmatrix}\mathbf{L}_{r}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}\end{bmatrix}\begin{bmatrix}\mathbf{i}_{r}^{(n)}\\ \mathbf{\varphi}_{r}^{(n)}\end{bmatrix}+\begin{bmatrix}\Delta t\mathbf{e}_{s,r}^{(n+1)}\\ \mathbf{0}\end{bmatrix}$, (24)
where $\mathbf{e}_{s,r}^{(n+1)}=\mathbf{e}_{s,r}(t_{n+1})$, and $\mathbf{R}_{r}$ again assembled from the DEIM-selected rows. This approximation linearizes the system in the unknowns, so advancing the state reduces to solving a linear system—no Newton iterations or Jacobian evaluations are required. However, care must be taken: if the problem is strongly nonlinear, this approximation may introduce large errors, since even a small error in the current can significantly affect $\mathbf{R}_{r}$, and hence the computed next-state current. For the highly nonlinear problem addressed in this paper, this approach proves infeasible, as demonstrated in Section 6.6.
@@ 5.1 Model Definition
Instead of approximating the nonlinear operator $\mathbf{R}(\mathbf{i})$ using POD-DEIM, we introduce a neural-network-based surrogate that operates entirely in the POD-reduced current space.
We consider the nonlinear function:
$\mathbf{R}_{r}:\mathbb{R}^{r_{i}}$ $\longrightarrow\mathbb{R}^{r_{i}\times r_{i}}$ (25)
$\mathbf{i}_{r}$ $\longmapsto\mathbf{R}_{r}(\mathbf{i}_{r}),$
that gives the resistance operator in the reduced space as a function of the reduced current-state. This function is modeled as a feed-forward neural network, denoted by $\mathbf{R}_{\psi}$, where $\psi$ represents the set of trainable parameters (weights and biases) of the network. The network consists of sequential layers that apply an affine transformation followed by a smooth, element-wise nonlinearity.

### v2

# A Structured Neural ODE Approach for Real-Time Evaluation of AC Losses in 3D Superconducting Tapes

@@ Abstract
Efficient modeling of High Temperature Superconductors (HTS) is crucial for real-time quench monitoring; however, full-order electromagnetic simulations remain prohibitively costly due to the strong nonlinearities. Conventional projection-based reduced-order modeling pipelines for nonlinear problems, such as Proper Orthogonal Decomposition (POD)Discrete Empirical Interpolation Method (DEIM), alleviate this cost but often require intrusive access to the Full Order Model (FOM) operators and a substantial number of interpolation points for hyperreduction. This work investigates reduced-order strategies for Integral Equation Method (IEM) of HTS systems. We present the first application of PODDEIM to IEM-based HTS models, and introduce a Structured Neural Ordinary Differential Equation (Neural ODE) approach that learns nonlinear dynamics directly in the reduced space. The benchmark results show that Neural ODE outperforms PODDEIM both in efficiency and accuracy, highlighting its potential for real-time simulations of superconductors.
@@ 1 Introduction
Although HTS offer the major advantage of operating at significantly higher temperatures than conventional Low Temperature Superconductors (LTS), monitoring the quench phenomenon remains a critical challenge. This difficulty arises from the relatively slow reduction of the critical current $I_{c}$ with increasing temperature [20, 33]. The sudden transition from the superconducting state to the normal resistive state generates localized hot spots through the Joule effect, which, if not properly controlled, can lead to irreversible damage of the conductor. Reliable quench detection and mitigation are therefore essential for the safe and continuous operation of future fusion demonstrator power plants, such as the European DEMO [35].
@@
Modeling quench phenomena requires solving multiphysics problems that couple electromagnetics (EM), thermal behavior, and fluid dynamics [52]. While lumped-element models have been proposed, capturing the underlying physics with sufficient accuracy typically demands finite element analysis. For superconducting structures, the EM problem can be addressed using several numerical formulations, most notably the Finite Element Method (FEM) with magnetic field ($H$) formulations [43], or the TTAA approach [19]. An alternative is IEM, which is traditionally used in open-domain problems such as antenna radiation, but has also proven effective in computing AC losses in superconductors [30, 31].
@@
Due to the strong nonlinearities in the resistivity model, both FEM and IEM simulations require substantial computational resources to capture the transient, nonlinear dynamics of realistic 3D superconducting structures. This high computational cost renders them impractical for real-time monitoring. To address this limitation, several studies have explored the use of surrogate models to accelerate the EM solution, particularly in the context of estimating AC loss [18, 54]. A first attempt to apply POD to HTS systems within the IEM framework was reported in [45]. However, in this work, no hyperreduction strategy was adopted, as the authors argued that the sparsity of the operator $\mathbf{R}$ kept the assembly cost manageable. This is true for relatively small problems, but as the mesh dimension increases, the assembly quickly becomes considerably more demanding. In such cases, Empirical Interpolation Method (EIM) [3] and its discrete counterpart [10] can be applied to alleviate this cost, and these techniques have already been successfully used in electromagnetic problems [17]. Nevertheless, DEIM suffers from two critical drawbacks in the present context: the strong nonlinearity and spatial locality of superconducting material laws require a large number of interpolation points, which reduces efficiency, and its intrusive nature demands full access to the high-fidelity model assembly routines, which limits practical applicability. Furthermore, since the governing equations are highly nonlinear, the reduced-order model must still be solved with an implicit time-integration scheme, which entails a large number of Newton iterations and repeated Jacobian evaluations. As a result, even with DEIM, the simulation cost remains significant.
@@
These challenges have motivated a growing interest in Machine Learning (ML)-based modeling strategies, which are typically non-intrusive and data-driven. Here, we provide only a brief overview of the main approaches, without aiming to be exhaustive; several recent reviews give a comprehensive account of this rapidly evolving area [23, 5, 7].
@@
Neural Operators such as DeepONet [29], Fourier Neural Operator [26], and graph neural operator variants [2] aim to learn mappings from function spaces, rather than between finite-dimensional vectors. This formulation enables them to generalize across families of input functions, achieving mesh-independent predictions and fast inference once trained. Their flexibility and non-intrusivity make them particularly attractive for parametric problems and large-scale simulations. However, they also present drawbacks: training requires substantial data, the networks can be very large, and their ability to handle complex geometries is limited. Graph-based neural operators alleviate this to some extent, but at the price of predicting the full-order field directly, which implies very large network outputs and nontrivial computational costs.
@@
Physics-Informed Neural Networks (PINNs) [39] incorporate the residual of the physical equation into the training loss, enabling the solution of forward and inverse problems even when data are scarce. PINNs are especially powerful for inverse problems and for settings where measurements are sparse. However, they tend to struggle with complex geometries, and their formulation is not well-suited for input-driven or control problems: a PINN is typically trained for a fixed set of inputs or excitations, and adapting it to new ones generally requires retraining the network. Additionally, their implementation requires formulating the governing equations as differentiable residuals, which can be cumbersome when the full-order model is complex or not available in closed form.
@@
Latent Dynamics modeling frameworks [6, 13] take yet another route. The central idea is that the solution trajectories of high-dimensional problems evolve on a low-dimensional manifold, which can be identified either with linear projections (e.g. POD) or with nonlinear encoders (e.g., autoencoders [32, 24]). Once the manifold is fixed, the latent dynamics can be learned in different ways. Neural ODEs [11] model continuous-time dynamics by learning the time derivative as a neural network. Sparse Identification of Nonlinear Dynamics (SINDy) [8] also parametrizes the derivative in continuous time, but restricts it to a sparse linear combination of predefined candidate functions, leading to interpretable governing equations. More recent developments further combine physics-based structure and data-driven learning for parametric reduced-order modeling [36], and explore advanced nonlinear manifold-learning strategies such as probabilistic manifold decomposition [16], as well as autoencoder-based ROMs enhanced through self-attention mechanisms [14].
@@
In this work, we build on the latent dynamics paradigm and propose a novel neural architecture for nonlinear electromagnetic modeling that operates directly in the reduced-order space. The starting point is a discretized integral equation model of HTS conductors, which is compressed through POD-based dimensionality reduction. Instead of relying on intrusive hyper-reduction techniques such as DEIM, we design a Structured Neural ODE that learns only the nonlinear component of the reduced dynamics, while the linear part is inherited directly from the full-order model. This idea of embedding problem structure into Neural ODEs architectures has also been explored in recent works on Structured Neural ODEs [27, 41, 28], albeit motivated primarily by stiffness and time-integration considerations. In this way, physical knowledge is embedded into the reduced-order evolution, ensuring parsimony, interpretability, and efficiency. The approach is validated on a HTS CORC® cable case study and benchmarked against the standard PODDEIM method, demonstrating the potential of physics-informed latent dynamics modeling for challenging electromagnetic simulation. The results demonstrate improved accuracy and significantly lower computational cost, highlighting the potential of the proposed strategy for real-time, industrially relevant superconducting applications.
@@
  • We introduce and validate a Structured Neural ODE approach that directly learns the nonlinear dynamics in the reduced-order space, offering a less-intrusive alternative to traditional hyperreduction techniques.
@@ 2 Model Description
Figure 1: Geometrical model of HTS tapes under analysis (units in $m$), and 2D surface mesh.
@@
We investigate the current density distribution induced by an external time-varying magnetic field. Specifically, the magnetic field is assumed to be oriented in the $y$-direction. This external field is modeled as:
@@
where $B_{0}$ is the magnetic field amplitude , and $f=50$ Hz is the frequency. According to Faraday’s law of induction, this time-varying magnetic field induces an electric field:
@@ 3 $J-\varphi_{e}$ Formulation
Although this formulation leads to non-sparse matrices, it is advantageous because only the regions carrying current–namely, the superconducting domains–need to be meshed. This avoids the need to mesh the surrounding air, unlike the $H$ and $H-\varphi_{e}$ formulations, which require solving for air elements as well. As a result, the overall model size is significantly reduced.
@@ 4.1 Structure-Preserving POD
    Using FOM simulations, snapshots of the current and potential are collected:
@@
    where $n_{e}$, $n_{f}$, and $n_{s}$ are the number of edges, faces, and snapshot instances, respectively. Each column of the snapshot matrices corresponds to the system state at a given time instant, and the variability across snapshots arises from both temporal evolution and changes in the simulation setup, such as different inputs or initial conditions, depending on the specific application.
@@
    $\mathbf{V}_{i,r}\in\mathbb{R}^{n_{e}\times r_{i}},\quad\mathbf{V}_{\varphi,r}\in\mathbb{R}^{n_{f}\times r_{\varphi}}.$ (14)
@@ 4.3 POD-DEIM ROM Simulation
To this end, we approximate the time derivatives over $t_{n},t_{n+1}]$ using a first-order backward finite difference, namely
@@
$\begin{bmatrix}\Delta t\mathbf{R}_{deim}\big(\mathbf{i}_{r}^{(n+1)}\big)+\mathbf{L}_{r}&\Delta t\mathbf{G}_{r}^{T}\\ \Delta t\mathbf{G}_{r}&\mathbf{0}\end{bmatrix}\begin{bmatrix}\mathbf{i}_{r}^{(n+1)}\\ \mathbf{\varphi}_{r}^{(n+1)}\end{bmatrix}=\begin{bmatrix}\mathbf{L}_{r}\mathbf{0}\\ \mathbf{0}&\mathbf{0}\end{bmatrix}\begin{bmatrix}\mathbf{i}_{r}^{(n)}\\ \mathbf{\varphi}_{r}^{(n)}\end{bmatrix}+\begin{bmatrix}\Delta t\mathbf{e}_{s,r}^{(n+1)}\\ \mathbf{0}\end{bmatrix}$, (23)
with $\mathbf{R}_{deim}$ is assembled using only the rows selected by the DEIM procedure. This discretization leads to a nonlinear system, since $\mathbf{R}_{deim}$ depends nonlinearly on $\mathbf{i}_{r}^{(n+1)}$. Solving it requires a Newton solver, which in turn involves repeated Jacobian evaluations. This process is computationally expensive. To reduce the cost, we can, in principle, introduce an approximation: we evaluate $\mathbf{R}_{deim}$ at the current from the previous time step $\mathbf{i}_{r}^{(n)}$, rather than at the unknown current $\mathbf{i}_{r}^{(n+1)}$:
$\begin{bmatrix}\Delta t\mathbf{R}_{deim}\big(\mathbf{i}_{r}^{(n)}\big)+\mathbf{L}_{r}&\Delta t\mathbf{G}_{r}^{T}\\ \Delta t\mathbf{G}_{r}&\mathbf{0}\end{bmatrix}\begin{bmatrix}\mathbf{i}_{r}^{(n+1)}\\ \mathbf{\varphi}_{r}^{(n+1)}\end{bmatrix}=\begin{bmatrix}\mathbf{L}_{r}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}\end{bmatrix}\begin{bmatrix}\mathbf{i}_{r}^{(n)}\\ \mathbf{\varphi}_{r}^{(n)}\end{bmatrix}+\begin{bmatrix}\Delta t\mathbf{e}_{s,r}^{(n+1)}\\ \mathbf{0}\end{bmatrix}$, (24)
where $\mathbf{e}_{s,r}^{(n+1)}=\mathbf{e}_{s,r}(t_{n+1})$, and $\mathbf{R}_{deim}$ again assembled from the DEIM-selected rows. This approximation linearizes the system in the unknowns, so advancing the state reduces to solving a linear system–no Newton iterations or Jacobian evaluations are required. However, care must be taken: if the problem is strongly nonlinear, this approximation may introduce large errors, since even a small error in the current can significantly affect $\mathbf{R}_{deim}$, and hence the computed next-state current. For the highly nonlinear problem addressed in this paper, this approach proves infeasible, as demonstrated in Section 6.7.
@@ 5.1 Model Definition
We aim at constructing a non-intrusive ROM that avoids hyperreduction of the nonlinear operator. Starting from the POD-reduced formulation, the reduced dynamics can be written in the form
$\begin{bmatrix}\mathbf{R}_{r}(\mathbf{i_{r}})&\mathbf{G}_{r}^{T}\\ \mathbf{G}_{r}&\mathbf{0}\end{bmatrix}\begin{bmatrix}\mathbf{i}_{r}\\ \mathbf{\varphi}_{r}\end{bmatrix}+\begin{bmatrix}\mathbf{L}_{r}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}\end{bmatrix}$ $\frac{d}{dt}\begin{bmatrix}\mathbf{i}_{r}\\ \mathbf{\varphi}_{r}\end{bmatrix}=\begin{bmatrix}\mathbf{e}_{s,r}(t)\\ \mathbf{0}\end{bmatrix}$. (25)
where the linear operators $\mathbf{L}_{r}$ and $\mathbf{G}_{r}$ are inherited from POD projection of the FOM, while the only remaining nonlinear contribution is the reduced resistance operator $\mathbf{R}_{r}(\mathbf{i}_{r})$:
$\mathbf{R}_{r}:\mathbb{R}^{r_{i}}$ $\longrightarrow\mathbb{R}^{r_{i}\times r_{i}},$ (26)
$\mathbf{i}_{r}$ $\longmapsto\mathbf{R}_{r}(\mathbf{i}_{r}).$
Instead of approximating $\mathbf{R}_{r}(\mathbf{i}_{r})$ via POD-DEIM, we introduce a neural-network surrogate that operates entirely in the reduced current space. Specifically, we approximate the reduced resistance operator with a feed-forward neural network
$\widehat{\mathbf{R}}_{r}(\mathbf{i}_{r};\psi)=\mathrm{NN}_{\psi}(\mathbf{i}_{r}),\qquad\mathrm{NN}_{\psi}:\mathbb{R}^{r_{i}}\rightarrow\mathbb{R}^{r_{i}\times r_{i}},$ (27)
where $\psi$ denotes the trainable parameters (weights and biases). For ease of notation, in the remainder of the manuscript, we denote this neural surrogate simply by $\mathbf{R}_{\psi}(\mathbf{i}_{r})$. The network takes as input the reduced current degrees of freedom $\mathbf{i}_{r}$ and outputs the entries of the reduced resistance operator. Substituting $\mathbf{R}_{\psi}(\mathbf{i}_{r})$ into (25) yields a reduced dynamical system parametrized by $\psi$; in this sense, the neural network defines a learnable vector field and the resulting model falls within the Neural ODE paradigm.